Complete.
14 ft2 = _______ in.2

Answer by rapaljer(4671)

In 1 sq ft, there are 12*12= 144 sq in.
14 sq ft = 14*144= 2016 sq in, or 2016 in^2.

Dr. Robert J. Rapalje, Retired